Resnick Inc. recently hired you as a financial analyst. Your first task is to assess the company's funding and to determine the firm's cost of capital. In your research you find that the company's target capital structure is 40% debt, 50% common equity and 10% preferred equity and that the company has no plans for issuing new stock so all common equity comes from retained earnings. Resnick's common stock has a current price of $60, a beta of 1.4, an upcoming dividend of $3 and a constant growth rate of 5%. Resnick's outstanding bonds have 15 years to maturity, a coupon rate of 6% paid annually, par value of $1,000, and they are currently trading for $758.15. The company is in the 30% tax bracket. The company also has preferred stock with par value of $100 and an 8% preferred dividend. Each preferred share currently trades for $80. The stock is currently trading for $60. If the risk-free rate is 2% and the market risk premium is 7%, what should be Resnick's cost of capital?
Answer #1
Weight of equity = E/A
Weight of equity =
W(E)=0.5
Weight of debt = D/A
Weight of debt = 0.4
W(D)=0.4
Weight of preferred equity =1-D/A-E/A
Weight of preferred equity = =1-0.4 - 0.5
W(PE)=0.1
Cost of equity
As per DDM
Price = Dividend in 1 year/(cost of equity - growth rate)
60 = 3/ (Cost of equity - 0.05)
Cost of equity% = 10
Cost of debt
                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =15
758.15 =∑ [(6*1000/100)/(1 + YTM/100)^k]     +   1000/(1 + YTM/100)^15
                   k=1
YTM = 9.0004424436
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 9.0004424436*(1-0.3)
= 6.30030971052
cost of preferred equity
cost of preferred equity = Preferred dividend/price*100
cost of preferred equity = 8/80*100
=10
WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=6.3*0.4+10*0.5+10*0.1
WACC =8.52%